# Learning Distances from Data with Normalizing Flows and Score Matching

### v1

@@ Abstract
Density-based distances (DBDs) offer an elegant solution to the problem of metric learning. By defining a Riemannian metric which increases with decreasing probability density, shortest paths naturally follow the data manifold and points are clustered according to the modes of the data. We show that existing methods to estimate Fermat distances, a particular choice of DBD, suffer from poor convergence in both low and high dimensions due to i) inaccurate density estimates and ii) reliance on graph-based paths which are increasingly rough in high dimensions. To address these issues, we propose learning the densities using a normalizing flow, a generative model with tractable density estimation, and employing a smooth relaxation method using a score model initialized from a graph-based proposal. Additionally, we introduce a dimension-adapted Fermat distance that exhibits more intuitive behavior when scaled to high dimensions and offers better numerical properties. Our work paves the way for practical use of density-based distances, especially in high-dimensional spaces.
@@ 1 Introduction
Metric learning aims to determine a distance metric that accurately measures the similarity or dissimilarity between data points. Typically, data points are mapped into a space where simple metrics like Euclidean distance can be applied. Although this method is computationally efficient, it is limited by the configurations possible in Euclidean space, where arbitrary distance relationships between sets of points cannot be represented.
@@
A more flexible but computationally expensive approach involves defining a Riemannian metric in the data space and solving for geodesic distances. This allows for arbitrary distance relationships, but it requires choosing an appropriate Riemannian metric. Fermat distances (Groisman et al., 2022), which are a type of density-based distance (Bousquet et al., 2003), offer an elegant solution: the metric should be conformal (a multiple of the identity matrix) and inversely proportional to the probability density. In this way, geodesics pass through high-density regions of the data, respecting any inherent manifold structure. This approach is consistent with Fermat’s principle of least time in optics, where the inverse density functions like a refractive index.
To address the computational challenge of computing geodesics with this metric, previous work has approximated trajectories using shortest paths in a graph of nearest neighbors (Bijral et al., 2012). The edge weights are approximated by raising Euclidean distances to an appropriate power, with theoretical analysis supporting the consistency of this approximation. However, in practice, we observe that convergence quickly plateaus, showing no significant improvement with increasing sample size.
@@
By employing more accurate edge weights obtained through normalizing flows, we achieve much faster convergence. Additionally, previous approaches did not account for the limitations of graph-based methods in higher dimensions, where paths become increasingly rough due to the curse of dimensionality. To address this, we use a relaxation scheme to smooth trajectories and maintain satisfactory convergence. We find that training models via score matching is more effective for this purpose than using normalizing flows.
@@
In summary, our contributions are threefold: First, we introduce the use of normalizing flows to obtain more accurate edge weights for estimating Fermat distances, significantly improving convergence rates. Second, we address the limitations of graph-based methods in high-dimensional spaces by implementing a smooth relaxation scheme using score matching, which better maintains the continuity and smoothness of the trajectories. Finally, we propose a dimension-adapted Fermat distance that scales intuitively with increasing dimensionality and exhibits superior numerical properties, paving the way for practical applications of density-based distances in complex, high-dimensional datasets. Through these innovations, we demonstrate the practicality and effectiveness of our approach, opening up a new possibility for metric learning using density-based distances in high-dimensional spaces.
Figure 1: Ground truth and estimated geodesics on a 2-dimensional spiral dataset. The green line shows a ground truth geodesic, solved via relaxation on the true density (Algorithm 1). The dashed orange line is an estimate for the geodesic, found as the shortest path in a nearest neighbor graph with edge weights calculated from a normalizing flow density (Algorithm 2).
2 Related work
Bousquet et al. (2003) introduces the idea of changing the geometry of the data space by use of a conformal density-based metric in order to get more informative distance functions. However they do not attempt to calculate geodesics or distances implied by this geometry, instead linking it to a modified SVM algorithm used for semi-supervised learning. Sajama and Orlitsky (2005) also uses a density-based metric for semi-supervised learning. They estimate densities using kernel density estimation and construct a graph of nearest neighbor distances from which they compute geodesics.
Bijral et al. (2012) continues this line of work by approximating the density as inversely proportional to a power of the Euclidean distance between close neighbors. Hwang et al. (2016) gives convergence guarantees for the power-weighted shortest path distances to geodesic distances under the density-dependent metric. Several other papers give similar results, and some practical implementations (Chu et al., 2020; Little et al., 2022; Groisman et al., 2022; Moscovich et al., 2017; Alamgir and Von Luxburg, 2012; Mckenzie and Damelin, 2019; Little et al., 2020). Terms used are power-weighted shortest path distances (PWSPD) (Little et al., 2022) and Fermat distances (Groisman et al., 2022; Trillos et al., 2023).
One crucial contribution of our work is to check that proposed methods actually converge in practice, regardless of theoretical guarantees. Previous work does not compare shortest paths to ground truth geodesics, even for a uniform distribution, where geodesics are simply straight lines. We address this major shortcoming by computing the ground truth geodesics for problems with known density and find that existing methods show very poor convergence.
@@ 3.1 Riemannian geometry
Riemannian geometry offers powerful tools for analyzing curved spaces, which can be particularly useful for understanding and optimizing over non-Euclidean spaces such as manifolds. A fundamental object in this field is the metric tensor, where different choices of metric can lead to vastly different geometric structures, such as positive curvature (e.g., a sphere which closes in on itself) or negative curvature (e.g., hyperbolic space which expands exponentially quickly), as well as more complicated mixtures of the two.
More formally, a Riemannian manifold $\mathcal{M}$ is a smooth $n$-dimensional space that locally resembles $\mathbb{R}^{n}$. At each point $p$ on the manifold, there is an associated tangent space $\mathcal{T}_{p}$, which is a linear approximation of the manifold at $p$. The metric tensor $g$ is a smoothly varying positive-definite bilinear form on the tangent space. It defines an inner product in the tangent space, allowing us to measure lengths and angles.
[…]
The distance between two points on the manifold is defined as the minimal length of a curve connecting the points. The curves that minimize this length are known as geodesics, which are the generalization of straight lines in Euclidean space to curved spaces.
@@ 3.2 Fermat distances
Density-based distances (DBDs) provide a way to define distances between points in a distribution by utilizing the underlying probability density function. This approach can be particularly useful for tasks such as clustering and pathfinding in data analysis and machine learning. If designed well, a DBD can capture desirable properties such as:
@@
$g(u,v)=\frac{\langle u,v\rangle}{f(p)^{2}},$ (1)
@@
3.2.1 Fermat distances which scale with dimension
Figure 2: Geodesics of the standard normal distribution, projected into 2d, taken between two orthogonal points at distance $\sqrt{D}$ from the origin. The paths are scaled by $\sqrt{D}$ to enable comparison. (Left) Unscaled metric (same temperature for all dimensions) leads to sharply curved geodesics in high dimensions. (Right) Scaled metric (temperature equal to dimension) leads to consistent geodesics in all dimensions (note that the lines are overlapping).
If $\beta=1$, this leads to unintuitive behavior in higher dimensions. Consider the standard normal distribution in $D$ dimensions. Due to rotational symmetry, the geodesic connecting $x_{1}$ and $x_{2}$ will lie in the plane spanned by these two points and the origin. Without loss of generality, suppose this plane aligns with the first two axis directions. The density restricted to this plane is the same as in the two-dimensional case, up to a constant factor. Therefore, geodesics in the $D$-dimensional standard normal distribution are the same as in the two-dimensional standard normal distribution, once we account for the rotation into the plane of $x_{1}$ and $x_{2}$.
However, despite this similarity, the typical distance from the origin of points sampled from the $D$-dimensional standard normal distribution is much larger, with a mean of $\sqrt{D}$. As a result, geodesics will start and end in very low-density regions relative to the two-dimensional standard normal. This means that most of the trajectory is spent moving towards and then outwards from the high-density region near the origin.
To address this issue, if we scale the temperature linearly with dimension ($\beta=1/D$), the results become consistent with the two-dimensional case (see Fig. 2). Scaling the temperature with dimension also increases numerical stability. This is because the effect of the probability density is reduced, leading to more Euclidean-like geodesics. This effect is clearly visible in Fig. 2, where the geodesics are much closer to straight lines in the scaled case.
@@ 3.3 Geodesic equation
From the metric, we can compute the geodesic equation. This is the second-order differential equation which all paths of minimum length (geodesics) must satisfy. In our case, we have the relatively simple solution for a trajectory $\gamma(t)$:
@@
The geodesic equation has the property that the speed is constant with respect to the metric, meaning $\sqrt{g(\dot{\gamma},\dot{\gamma})}$ is constant. In other words, the geodesic distance traveled in a given time interval is the same at all points of the curve. If $\gamma$ passes from a high-probability region to a low-probability one, the distances in the low-probability region could be orders of magnitude larger than in the high-probability region. This makes numerically solving the equations challenging due to the vastly different Euclidean velocities along the trajectory.
It is convenient to reparameterize the equation such that the Euclidean velocity is constant instead, as this is much easier to handle numerically. This is especially useful for solving via relaxation, where having equal-length segments leads to much more stable behavior. Making this change results in a very similar equation (removing only the factor of 2 in the second term):
@@
See Section A.1 for derivations of these results.
4 Methods
4.1 Ground truth distances through relaxation
A relaxation scheme allows one to solve a differential equation constrained by boundary conditions by iteratively updating the solution such that it better obeys the differential equation. We divide the interval $[0,1]$ into equal time steps of size $h=1/n$ and write $\varphi_{i}=\varphi(ih)$. Initialize the values of the curve $\varphi$ at $t=ih$ for $i=0,\dots,n$, with the constraint that $\varphi_{0}$ and $\varphi_{n}$ are the chosen endpoints. The initial curve could be the straight line connecting the end points, or a more informative guess if additional information is available.
Equation 4 leads to the following relaxation scheme. Define
$v_{i}=\frac{\varphi_{i+1}-\varphi_{i-1}}{2}$ (5)
then update with
$\varphi^{\prime}_{i}=\frac{\varphi_{i+1}+\varphi_{i-1}}{2}+\frac{\beta}{2}\left(s(\varphi_{i})\|v_{i}\|^{2}-(s(\varphi_{i})\cdot v_{i})v_{i}\right)$ (6)
See Section A.2 for a derivation.
4.1.1 Implementation
In practice, some additional features are needed to ensure stable convergence of the relaxation scheme, especially when the score function is not particularly smooth. Firstly, we found that oscillatory features commonly arise. To address this, we use smoothing after the update, returning a value which is interpolated between the value at the node and the average of its neighbors. Secondly, for rougher score functions we use a step size between 0 and 1, where 0 corresponds to no update and 1 corresponds to the ordinary relaxation step. See Algorithm 1.
Algorithm 1 One step of relaxation. $\varphi=\{\varphi_{i}\}_{i=0}^{n}$ is the current state of the trajectory, $s$ is the score function, $\beta$ the inverse temperature, $\gamma\in[0,1)$ the smoothing parameter and $\alpha\in(0,1]$ the step size. This step should be iterated until convergence, which can be measured via the difference between $\varphi$ and $\varphi^{\prime}$.
function RelaxationStep($\varphi,s,\beta,\gamma,\alpha$)
     $v_{i}\leftarrow\frac{1}{2}(\varphi_{i+1}-\varphi_{i-1})$triangleright $i=1,\dots,n-1$
     $\varphi^{\prime}_{i}\leftarrow\frac{1}{2}(\varphi_{i+1}+\varphi_{i-1})+\frac{\beta}{2}\left(s(\varphi_{i})\|v_{i}\|^{2}-(s(\varphi_{i})\cdot v_{i})v_{i}\right)$
     $\varphi^{\prime}_{i}\leftarrow(1-\gamma)\varphi^{\prime}_{i}+\frac{\gamma}{2}(\varphi^{\prime}_{i+1}+\varphi^{\prime}_{i-1})$
     $\varphi^{\prime}_{i}\leftarrow\alpha\varphi^{\prime}_{i}+(1-\alpha)\varphi_{i}$
     return $\varphi^{\prime}$
end function
@@ 4.2 Weighted graph algorithm based on Euclidean distance
If we have a large number of samples from $p(x)$, we can approximate the density between close neighbors $x_{1}$ and $x_{2}$ as constant and hence the shortest path between them as a straight line. The density is roughly proportional to the inverse of the Euclidean distance, raised to the power $d$, where $d$ is the intrinsic dimension of the distribution (Bijral et al., 2012):
@@
Bijral et al. (2012) therefore proposes to approximate shortest paths between more distant points by constructing a $k$-nearest neighbors graph of the samples with edge weights given by $\|x_{1}-x_{2}\|^{\beta d+1}$ and applying Dijkstra’s algorithm.
@@
While this approach is appealing in its simplicity, can be easily implemented using standard libraries, and comes with theoretical guarantees, the density estimate is too poor to provide accurate estimates of Fermat distances in practice. In the experimental section 5.1.1 we show that paths computed with this approach do not converge to the ground truth geodesics as sample size increases, even in very simple 2-dimensional distributions such as the standard normal.
We are unsure why we observe such a large deviation from the theoretical prediction. It is possible that the unknown constant $\theta_{0}$ is very small, leading to extremely slow convergence, despite the exponential function. We leave investigation into this point to future work.

### v2

@@ Abstract
Density-based distances (DBDs) provide a principled approach to metric learning by defining distances in terms of the underlying data distribution. By employing a Riemannian metric that increases in regions of low probability density, shortest paths naturally follow the data manifold. Fermat distances, a specific type of DBD, have attractive properties, but existing estimators based on nearest neighbor graphs suffer from poor convergence due to inaccurate density estimates. Moreover, graph-based methods scale poorly to high dimensions, as the proposed geodesics are often insufficiently smooth. We address these challenges in two key ways. First, we learn densities using normalizing flows. Second, we refine geodesics through relaxation, guided by a learned score model. Additionally, we introduce a dimension-adapted Fermat distance that scales intuitively to high dimensions and improves numerical stability. Our work paves the way for the practical use of density-based distances, especially in high-dimensional spaces.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Our method learns distance metrics that respect the underlying data distribution. Here, we show shortest paths between two points in a complex Gaussian mixture distribution. The ground truth geodesic (dashed red) and our graph-based approximation (solid blue) both naturally follow high-density regions where data samples (green dots) are concentrated, rather than taking a direct Euclidean path. This behavior is desirable for many machine learning tasks, as it captures the intrinsic structure of the data manifold.
Metric learning is fundamental to many machine learning tasks, from clustering to classification. The goal is to determine a distance metric that accurately measures the similarity or dissimilarity between data points. While traditional approaches map data points into spaces where Euclidean distance can be directly applied, these methods are inherently limited by the geometric constraints of Euclidean space, making it impossible to represent arbitrary distance relationships between sets of points.
@@
A more flexible but computationally expensive approach involves defining a Riemannian metric in the data space and solving for geodesic distances. This allows for arbitrary distance relationships, but it requires choosing an appropriate Riemannian metric. Fermat distances (Groisman et al., 2022), which are a type of density-based distance (Bousquet et al., 2003), offer an elegant solution: the metric should be conformal (a multiple of the identity matrix) and inversely proportional to the probability density. Intuitively, this means that distances are stretched in regions where data is sparse and compressed where data is dense. In this way, geodesics pass through high-density regions of the data, respecting any inherent manifold structure (see Figure 1). This approach is consistent with Fermat’s principle of least time in optics, where the inverse density functions like a refractive index.
This density-based approach to metric learning has evolved significantly over the past two decades. The foundational work of Bousquet et al. (2003) introduced the concept of conformal density-based metrics, though their focus was on semi-supervised learning rather than computing geodesics. Sajama & Orlitsky (2005) took the first steps toward practical implementation by using kernel density estimation and nearest-neighbor graphs for geodesic computation. A breakthrough came from Bijral et al. (2011), who approximated the density as inversely proportional to a power of the Euclidean distance between close neighbors. This elegant approximation sparked numerous theoretical extensions and practical implementations (Chu et al., 2020; Little et al., 2022; Groisman et al., 2022; Moscovich et al., 2017; Alamgir & Von Luxburg, 2012; Mckenzie & Damelin, 2019; Little et al., 2020), becoming known as power-weighted shortest path distances (PWSPD) (Little et al., 2022) or Fermat distances (Groisman et al., 2022; Trillos et al., 2023). While Hwang et al. (2016) provided theoretical convergence guarantees for these methods, notably absent from all previous work is any comparison to ground truth distances, even in simple cases where they can be computed exactly. Our empirical analysis reveals that existing approaches exhibit poor convergence in practice, highlighting a significant gap between theoretical guarantees and practical performance.
Other notable works which use density-based metrics to inform the learning of data manifolds include metric flow matching (Kapusniak et al., 2024) and the Riemannian autoencoder (Diepeveen et al., 2024), which uses a metric defined in terms of score functions.
To establish ground truth geodesics and distances, we develop a numerically stable relaxation method that directly solves the geodesic differential equations. By carefully handling the reparameterization of paths to maintain constant Euclidean velocity and employing appropriate numerical optimization techniques, we can compute accurate geodesics even in challenging cases with large density variations. This enables us, for the first time, to quantitatively evaluate the accuracy of different approximation methods by comparing them to true geodesics in distributions with known densities, such as mixtures of Gaussians.
@@
With this ground truth framework in place, we develop two key improvements to existing methods. First, by employing more accurate edge weights obtained through normalizing flows, we achieve much faster convergence in graph-based approximations. Second, we address the limitations of these graph-based methods in higher dimensions, where paths become increasingly rough due to the curse of dimensionality, by using a relaxation scheme to smooth trajectories. Interestingly, we find that training models via score matching is more effective than using normalizing flows for this purpose, revealing a trade-off between the accuracy of learned log densities and their gradients (scores).
@@
In summary, our contributions are fourfold: First, we introduce a numerically stable relaxation method to compute ground truth geodesics and distances. Second, we introduce the use of normalizing flows to obtain more accurate edge weights for estimating Fermat distances, achieving significantly faster convergence rates than previous methods. Third, we overcome the limitations of graph-based methods in high-dimensional spaces through a novel smooth relaxation scheme using score matching. Finally, we propose a Fermat distance that naturally scales with dimensionality, making density-based distances practical for real-world, high-dimensional applications. Our approach bridges the gap between theoretical elegance and practical applicability in metric learning.
@@ 2.1 Riemannian Geometry
Riemannian geometry provides the mathematical foundation for understanding curved spaces and measuring distances within them. While Euclidean geometry suffices for flat spaces, many real-world datasets lie on curved manifolds where Euclidean distances can be misleading. Riemannian geometry offers powerful tools for analyzing such spaces through the metric tensor, which can encode different geometric structures ranging from positive curvature (like a sphere) to negative curvature (like hyperbolic space), or more complex combinations of both.
More formally, a Riemannian manifold $\mathcal{M}$ is a smooth $n$-dimensional space that locally resembles $\mathbb{R}^{n}$. In this work, we make the simplifying assumption that $\mathcal{M}=\mathbb{R}^{n}$. As a result the tangent space $\mathcal{T}_{p}$, which is a linear approximation of the manifold at $p$, is simply $\mathbb{R}^{n}$ for all $p$. The metric tensor $g_{p}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is a smoothly varying positive-definite bilinear form on the tangent space. It defines an inner product in the tangent space, allowing us to measure lengths and angles.
For a tangent vector $v\in\mathcal{T}_{p}$, the norm is defined as $\|v\|=\sqrt{g_{p}(v,v)}$, providing a notion of length in the tangent space. Just as we measure the length of a curve in Euclidean space by integrating its speed, the length of a smooth curve $\gamma(t)$ on the manifold is given by $\int\|\dot{\gamma}(t)\|\,\mathrm{d}t$, where $\dot{\gamma}(t)$ is the derivative of $\gamma(t)$ with respect to $t$. This integral accumulates the infinitesimal lengths along the curve, accounting for how the metric varies across the manifold.
The distance between two points on the manifold is defined as the minimal length of a curve connecting the points. The curves that minimize this length are known as geodesics, which are the generalization of straight lines in Euclidean space to curved spaces. These geodesics satisfy a second-order differential equation known as the geodesic equation (see Section A.1). In our context, these geodesics will represent optimal paths through the data space that respect the underlying probability density structure.
@@ 2.2 Fermat Distances
Having established the framework of Riemannian geometry, we now turn to density-based distances (DBDs), which provide a way to define distances between points sampled from a distribution by utilizing the underlying probability density function. This approach can be particularly useful for tasks such as clustering and pathfinding. If designed well, a DBD can capture desirable properties such as:
@@
$g_{x}(u,v)=\frac{\langle u,v\rangle}{f(p(x))^{2}},$ (1)
@@
This integral accumulates the Euclidean length element $\|\dot{\gamma}(t)\|$, weighted inversely by the density along the path. As a result, segments of the path that pass through low-density regions contribute more to the total length than those through high-density regions.
Fermat Distances Which Scale with Dimension
Figure 2: Geodesics of the standard normal distribution, projected into 2D, taken between two orthogonal points at distance $\sqrt{D}$ from the origin. The paths are scaled by $\sqrt{D}$ to enable comparison. Left: Unscaled metric (same temperature for all dimensions) leads to sharply curved geodesics in high dimensions. Right: Scaled metric (temperature equal to dimension) leads to consistent geodesics in all dimensions (note that the lines are overlapping).
While the Fermat distance framework is theoretically elegant, setting $\beta=1$ (as is common in previous work) leads to unintuitive behavior in higher dimensions. Consider the standard normal distribution in $D$ dimensions. Due to rotational symmetry, the geodesic connecting $x_{1}$ and $x_{2}$ will lie in the plane spanned by these two points and the origin. Without loss of generality, suppose this plane aligns with the first two axis directions. The density restricted to this plane is the same as in the two-dimensional case, up to a constant factor. Therefore, geodesics in the $D$-dimensional standard normal distribution are the same as in the two-dimensional standard normal distribution, once we account for the rotation into the plane of $x_{1}$ and $x_{2}$.
However, despite this geometric similarity, there’s a crucial scaling issue: the typical distance from the origin of points sampled from the $D$-dimensional standard normal distribution is much larger, with a mean of $\sqrt{D}$. As a result, geodesics will start and end in exponentially low-density regions relative to the two-dimensional standard normal. This forces most of the trajectory to make a dramatic detour towards and then away from the high-density region near the origin, creating increasingly extreme curved paths as dimension grows.
To address this issue, if we scale the temperature linearly with dimension ($\beta=1/D$), the results become consistent with the two-dimensional case (see Figure 2). In Section A.3, we explore this argument in more detail and generalize it to non-Gaussian distributions. This scaling has two key benefits: First, it ensures that geodesic behavior remains intuitive across dimensions, with similar qualitative properties regardless of the ambient dimension. Second, it significantly improves numerical stability by reducing the effect of the probability density, leading to more Euclidean-like geodesics. This effect is clearly visible in Figure 2, where the geodesics are much closer to straight lines in the scaled case.
@@ 2.3 Geodesic Equation
Figure 3: Comparison of geodesic parameterizations. Left: Constant metric speed leads to varying Euclidean step sizes, making numerical relaxation challenging. Right: Constant Euclidean speed enables uniform discretization and more stable numerics. Points along the paths indicate discretization steps. Note the slightly different solutions: the left-hand path is approximately 0.1% longer.
From the metric, we can compute the geodesic equation, which characterizes the paths of minimum length between points. Just as a straight line in Euclidean space can be described by a second-order differential equation ($\ddot{x}=0$), geodesics in our curved space satisfy a more complex equation (see Theorem A.2). For a geodesic trajectory $\gamma(t)$, we have:
@@
The geodesic equation has the property that the speed is constant with respect to the metric, meaning $\sqrt{g(\dot{\gamma},\dot{\gamma})}$ is constant. While mathematically elegant, this creates numerical challenges: when $\gamma$ passes from a high-probability region to a low-probability one, the Euclidean velocities can differ by orders of magnitude. For example, if the probability density drops by a factor of 1000, the Euclidean velocity must increase by the same factor to maintain constant metric speed. This dramatic speed variation makes standard numerical integrators unstable and requires extremely small step sizes in low-density regions while potentially overshooting in high-density regions.
To address these numerical challenges, we reparameterize the equation to maintain constant Euclidean speed instead of constant metric speed (see Theorem A.3). This simplification is particularly valuable for numerical methods like relaxation, where uniform spacing of points along the trajectory leads to better stability. Figure 3 illustrates how this reparameterization enables uniform discretization compared to the varying step sizes required by constant metric speed. The reparameterized equation differs only slightly from the original (losing just the factor of 2 in the second term):
@@
We stress that, while Equation 3 is easily obtained from standard results on conformal metrics (e.g., Carroll (2004, Appendix G)), the reparameterized form in Equation 4 is, to our knowledge, a novel contribution.
@@ 2.4 Weighted Graph Algorithm Based on Euclidean Distance
Figure 4: Illustration of the power-weighted graph approximation. Left: Sample points (dots) and density (gradient) from a 2D standard normal. Center: k-nearest neighbor connections between points, with edge weights determined by cubic Euclidean distances. Right: A shortest path (solid blue) found by Dijkstra’s algorithm compared to the true geodesic (dashed red). Note how the discrete path approximates the continuous geodesic by following high-density regions, but is not very smooth.
A practical approach to computing Fermat distances is to discretize the problem using sample points. If we have a large number of samples from $p(x)$, we can approximate the density between close neighbors $x_{1}$ and $x_{2}$ as constant and hence the shortest path between them as a straight line. Following Bijral et al. (2011), the density is roughly proportional to the inverse of the Euclidean distance, raised to the power $d$, where $d$ is the intrinsic dimension of the distribution:
@@
This elegant relationship suggests a simple algorithmic approach: Bijral et al. (2011) proposes to approximate shortest paths between more distant points by constructing a $k$-nearest neighbors graph of the samples with edge weights given by $\|x_{1}-x_{2}\|^{\beta d+1}$ and applying Dijkstra’s algorithm. Figure 4 illustrates the method.
@@
While this approach is appealing in its simplicity, can be easily implemented using standard libraries, and comes with theoretical guarantees, the density estimate is too poor to provide accurate estimates of Fermat distances in practice. In the experimental section 4.1.1 we show that paths computed with this approach converge extremely slowly to the ground truth geodesics as sample size increases, even in very simple 2-dimensional distributions such as the standard normal.
This significant gap between theoretical guarantees and practical performance leads us to speculate that the unknown constant $\theta_{0}$ is very small, resulting in extremely slow convergence despite the exponential form of the bound. This observation motivates our development of more accurate density estimation methods in the following sections. We leave a detailed theoretical investigation of this convergence behavior to future work.